OpenStudy (anonymous):

what is the solution to 5x-3(2x+1)=25-3x

OpenStudy (anonymous):

2X=26 ===>? X=13

OpenStudy (anonymous):

5x - 6x - 3 = 25 - 3x -1x - 3 = 25 - 3x 2x - 3 = 25 2x = 28 x = 14